Just Hip Ltd. manufactures and sells stainless steel coffee mugs. Expected mug sales for Just Hip (in units) for the next four months are as follows:

September

October

November

December

Budgeted unit sales

30,000

32,500

34,200

36,800

Just Hip, Ltd. likes to maintain a finished goods inventory equal to 40% of the next month's estimated sales. How many mugs should Just Hip plan on producing during the month of October?
Answer #1

October

Budgeted unit sales

32,500

Add: Ending inventory required (34200*40%) 13680
Total units required 46180
Less: Beginning Inventory (32500*40%) 13000
Production required during the month of October 33,180
Answer #2

Answer : 33,180 Mugs

units Must Produced in October = Oct Sales + Desired Ending inventory- Beginning Inventory

Units Must Produce = 32500+(40%of 34200)- (32500*40%) = 32500+13680-13000 = 33,180 Answer